OpenStudy (anonymous):

Simplify the expression 1 over 4 to the negative fourth power. 256 16 negative 1 over 16 negative 1 over 256

OpenStudy (anonymous):

\[\frac{1}{4^{-4}}\] Is this it?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Great we know that \[x^{-a}=\frac{1}{x^a}\] so with this information we can then simplify our expression

OpenStudy (anonymous):

\[\frac{1}{4^{-4}}=\frac{1}{\frac{1}{4^4}}=4^4\]

OpenStudy (anonymous):

Can you simplify ? \[4^4\]

OpenStudy (anonymous):

the answer would be 256 right

OpenStudy (anonymous):

yes you are right!

OpenStudy (anonymous):

256

OpenStudy (anonymous):

thanks man for the help